Calculate the current price of a stock given the following information: A dividend has a growth rate of 0% for the first two periods. Beginning period 3 and continuing in perpetuity, the dividend grows at 4%. The required rate of return is 7%, Do-S2.00 The current price of a stock would be S63 per share 9-7a Corporate Valuation Model A firm has $20 million of debt and $30 million of preferred stock. It has an expected free cash flow (FCF) of $3 million at the end of period 1 and S20 million at the end of period 2. The period 3 end FCF is expected to be$10 million and is then expected to grow at a rate of 4% in perpetuity. The WACC is 1 1%. If there are 2 million shares outstanding, what is the value per share? 9-8 Preferred Stock A preferred stock sells for $88/share and retums an $8 dividend. What is the preferred stocks expected return?
Answer #1

HI,

As per HOMEWORKLIB POLICY, we would be doing only first question here using dividend discount model.

here Current dividend D0 = $2

perpetual growth rate till 2 years = 0%

so after 2 years dividend = $2*(1+0%) = $2

perpetual growth rate after 2 years g= 4%

required rate of return k = 7%

so as per dividend discount model share price = Do*(1+g)/(k-g)

so share price after 2 years = 2*(1+4%)/(7%-4%) = 2*1.04/0.03 = $69.333

and current share price = 69.333/(1+7%)^2 = $60.56

Thanks